OpenStudy (anonymous):

Write an equation that you could solve using these steps: Add 7 to both sides. Then, divide both sides by 4. Explain how you know your answer is right.

OpenStudy (anonymous):

Lets take the steps and do it backwards. x = n (some number), lets say 2 x = 2 4x = 8 4x -7 = 1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Lets do the steps: 4x -7 +7 = 1+7, 4x=8 4x/4 = 8/4 x = 2

OpenStudy (anonymous):

it asked you to add by 7, then divided by 4...so starting from the end, multiply by 4, and subtract 7